| Operation Summary for January 25, 2012 - January 27, 2012 |
|
Wednesday January 25 The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam. During the day shift, the MI gap-clearing kickers were brought back online at 8:49 AM. Linac personnel switched to the I- Source at 9:04 AM. Operations established beam to the Meson FTBF experiment T-1018 at 10 AM. Neutron Therapy Facility (NTF) personnel began tuning at 10:15 AM due to the different source. Operations loaded protons into the Recycler at 10:43 AM. At 11:24 AM, a Linac RF station (LRF3) tripped of; it reset. Booster personnel began a study at 12:47 PM. At 2:58 PM, a Linac Klystron RF station (KRF5) tripped off; it reset. During the evening shift, at 5:13 PM, a Booster vacuum pump (BIPRF15) tripped off and took down BRF3 and 4 with it; everything recovered and reset. Booster personnel completed their study at 5:14 PM. At 6:02 PM, BRF4 tripped off for the first of two times in five hours; it reset. Operators completed a MI phase tune at 6:21 PM. Operations disabled T-1018 beam at 10:01 PM. Thursday January 26 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, at 5:04 AM, BRF4 tripped off for the first of two times in 40 minutes; it reset. At 7:29 AM, NTF personnel began preparing for treating two patients later in the day. During the day shift, KRF6 tripped off at 8:31 AM and would not reset; a Linac person investigated and made adjustments. The NuMI beam permit tripped off at 10:16 AM on a RF gap fault; it reset. Operations established beam to T-1018 around 10:40 AM. During the evening shift, At 4:16 PM, BRF17 tripped off for the first of four times in about 30 minutes; Booster and RF personnel investigated and reported that the power amplifier was bad and there was modulator problem as well – this would need an access to repair, which they will investigate further on Friday. At 6:42 PM, B:IPRF15 tripped off and brought down BRF2 and 4; everything recovered and reset. At 7:57 PM. B:IPRF12 tripped off and took down BRF11 and 12; everything recovered and reset. Operations disabled T-1018 beam at 9:59 PM. At 10:24 PM, the MI abort gap clearing kickers tripped off and would not reset; Operators contacted MI and Kicker personnel and then turned the kickers off. Friday January 27 The midnight shift began with Operations sending beam to the NuMI and MiniBooNE. During the midnight shift, a Linac quadrupole magnet power supply (L:QPS107) tripped off; it reset. At 4:43 AM, Operators prepared Enclosure C for an access around 8 AM. Operations Coordinator (OpsCo) Plans for Today The OpsCo agrees with the plans of Booster to access and repair BRF17. Starting on Tuesday, 1/31/12, the Main Injector will go down for 7 to 14 days to replace a quadrupole magnet (Q403); more on how this shutdown will affect the rest of the accelerator complex later. |
|
More Information Comments and Suggestions |
| last modified 01/27/2012 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |